OpenStudy (anonymous):

A $1,600,000 building is depreciated linearly over 40 years. The scrap value will be $0. 1. Find the annual rate of depreciation. a. -$1,600,000 b. -$40,000 c. $40,000 d. $1,600,000 e. $1,560,000 2. Find an expression for the building value in the t th year of use (0 < t < 40).

OpenStudy (anonymous):

\[1,6000,000\div 40=?\]

OpenStudy (anonymous):

these questions all have a certain sameness to them is this math by boredom ?

OpenStudy (anonymous):

I solved that and got 40,000. Then I used V= c-r which is 1600000 - 40000t but all of my possible answers start with 40,000 or -40,000 so I don't understand ....

OpenStudy (anonymous):

oh i guess you have to put a minus sign in front

OpenStudy (anonymous):

is one of them \[y=-40,000 t+1,600,000\]?

OpenStudy (anonymous):

Not boredom I am studying linear models and then once I get this down I move onto solving linear programming problems. I am just having a hard time getting this down ...

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yes that is one of my answers but I don't understand why if its v= c - r why this one is switched to v= r-c

OpenStudy (anonymous):

or even if its v(t) = m(t) + b. that makes sense but in the examples my professor gives its all c - r(t)

jimthompson5910 (jim_thompson5910):

#1 change in value = (new value) - (old value) change in value = (0) - (1,600,000) change in value = -1,600,000 change in years = (new year) - (old year) change in years = (40) - (0) change in years = 40 annual rate of depreciation = (change in value)/(change in years) annual rate of depreciation = (-1,600,000)/(40) annual rate of depreciation = -40,000 the item is losing $40,000 in value every year ------------------------------------------------------- # 2 from #1, the slope is -40,000 the y-intercept is 1.6 million because this is the starting value m = -40,000 b = 1.6 million so the y=mx+b equation would be y = -40,000x + 1,600,000 x = time in years (0 < x < 40) y = value of the item